Grade 11Algebra

if |z+1/z|=1;prove that for maximum |z|, z is purely imaginary

Nishtha Gahlot
8 Years agoGrade 11

1 Answer

ragini
8 Years ago
to prove [z+1/z] = 1  
imaginary numbers are the squreroot of non positive real numbers ,since zero   is non positive number and is its own squareroot , zero can be considered as imaginary number 
let number be in equation 
if z = 0 
[0+1/0]= 1
1/0=1
1/0 is undefined value so it can be the maximum value for z.